# MMAC-Copilot: Multi-modal Agent Collaboration Operating System Copilot

### v2

@@ Abstract
Autonomous virtual agents are often limited by their singular mode of interaction with real-world environments, restricting their versatility. To address this, we propose the Multi-Modal Agent Collaboration framework (MMAC-Copilot), a framework utilizes the collective expertise of diverse agents to enhance interaction ability with operating systems. The framework introduces a team collaboration chain, enabling each participating agent to contribute insights based on their specific domain knowledge, effectively reducing the hallucination associated with knowledge domain gaps. To evaluate the performance of MMAC-Copilot, we conducted experiments using both the GAIA benchmark and our newly introduced Visual Interaction Benchmark (VIBench). VIBench focuses on non-API-interactable applications across various domains, including 3D gaming, recreation, and office scenarios. MMAC-Copilot achieved exceptional performance on GAIA, with an average improvement of 6.8% over existing leading systems. Furthermore, it demonstrated remarkable capability on VIBench, particularly in managing various methods of interaction within systems and applications. These results underscore MMAC-Copilot’s potential in advancing the field of autonomous virtual agents through its innovative approach to agent collaboration.
[…]
1888
@@ 1 Introduction
Figure 1: Illustrates the sequence of actions undertaken by MMAC-Copilot in response to the user’s request "Play <Love Story> on Spotify". (A) Represents the initial state where the "Planner" formulates preliminary plans. . (B) "Programmer" launches the application using a bash command. (C) "Viewer" click the search box and enter <Love Story>. (D) "Viewer" click the play button.
@@
Large Language Models (LLMs) [1, 2, 3, 12, 27, 37, 38] has demonstrated unprecedented capabilities to tackle complex tasks that require human-like reasoning, decision-making, and collaborative efforts [11, 16, 28, 29, 45]. Soon afterwards, Large Vision Models (LVMs) [7, 21, 23, 24, 25, 54] expand beyond traditional text-based processing by integrating a visual dimension into LLMs. The applications of LVMs have reached diverse domains, with one notable example being their pivotal usage in autonomous virtual agents [44].
Usually, autonomous virtual agents require a unified interaction interface and a high degree of generalization to adapt to the challenges posed by the numerous different applications within an operating system. Although previous works [35, 44, 51] that utilize scripts and the planning abilities of LLMs have demonstrated success in some applications there are still some challenges: First, the acquisition of real-world environmental information relies on a single text modality which lead to a limited range of applications. Second, the hallucinations caused by the knowledge domain gap could lead to incorrect planning by virtual agents, resulting in extremely severe consequences.
To engage with the initial challenge, we introduce MMAC-Copilot, a framework designed to promote information exchange and collaboration among multi-modal agents. Specifically, we created five key agents: Planner, an expert in making plans, strategic thinking and resource allocation,ensuring that the collaborative efforts of all agents are directed to achieve the target; Librarian, proficient in managing question-answering tasks and API in retrieving information; Programmer, specializing in coding tasks, particularly skilled at executing Bash commands and Python programs; Viewer, focused on understanding image content for visual information and execute click interaction; and Video Analyst, dedicated to analyzing video content and extracting key events. In addition, there is an auxiliary agents: the Mentor, supervising system operations and agent interactions to ensure smooth functionality and integrity. When receiving a user query, Planner formulates an initial plan based on the user query’s requirements. It then engages with other agents through a team collaboration chain, allowing for a dynamic exchange of information and insight. As the plan progresses, it is continuously adapted based on the insights of each specialized agent.
In resolving the second dilemma, we introduce the team collaboration chain, allowing participating agents to adapt the initial plan crafted based on their domain expertise. Refer to the scenario depicted in Figure 3, where the user request is "Open the Discord application and send Hi to Dylan Li". Planner firstly delineates five sub-tasks: "Open Discord", "Navigate to friends", "Search for Dylan Li", "Open Chat with Dylan Li", and "Send a Message", all of which are assigned to Programmer. Upon completion of the first subtask by the Programmer, Mentor observes that Discord has been successfully launched and notes that "Dylan is already present in the application." Planner, upon receiving this feedback, revises the initial plan to "Open chat with Dylan" and "Send the message", and then reassigns it to Programmer. In this instance, Planner overlooks the fact that programmer needs to specifically integrate Discord’s proprietary API to accomplish these tasks. Adapting to the proprietary interfaces of thousands of applications to achieve universality is clearly impractical. Furthermore, unforeseen scenarios such as pop-up notifications or advertisements may disrupt task execution. Programmer fails to execute and returns the message, "Sorry, I don’t know how to do this." Upon receiving this message, Planner reallocates the task to Viewer. Viewer decomposes the coarse-grained task into finer-grained tasks: "click on Dylan", "type Hi", and "return the message", which are then successfully executed. Additionally, each subtask completion is reviewed by Mentor, and upon approval, the process advances to the next round. The process concludes after all rounds are completed.
@@
We assess MMAC-Copilot using two benchmarks: GAIA [26] and our newly introduced Visual Interaction Benchmark (VIBench). GAIA is designed to test the abilities of general artificial intelligence assistants. The MMAC-Copilot performed in level-1, level-2, and level-3 tasks with scores of 45.16, 20.75, and 6.12, respectively. It outperforms previous methods by 6.8%. VIBench, specifically developed for this study, tests the framework’s performance in non-API-interactable applications across diverse domains such as 3D gaming, recreation, and office environments. The results exhibits outstanding adaptability, handling interfaces with a high degree of accuracy, enhancing interaction within digitally environments.
@@
  • We propose MMAC-Copilot, a collaborative framework involving specialized agents designed to navigate and perform general operating system tasks. MMAC-Copilot enhances the interaction capabilities of autonomous virtual agents with operating systems by leveraging multi-modality in processing to tasks. The framework consists of distinct modal agents: "Librarian" for question-answering and information retrieval, "Programmer" for executing code and bash commands, "Viewer" for interpreting visual information, and "Video Analyst" for analyzing video content. By leveraging the collaboration of modal agents, MMAC-copilot can handle a diverse range of tasks with enhanced accuracy.
  • To mitigate the incorrect planning issue caused by LLMs’ hallucination, we introduce the team collaboration chain which allows participating agents to adjust the initial plan based on their domain expertise into MMAC-Copilot. It has demonstrated adaptability in handling a variety of case.
  • We develop the Visual Interaction Benchmark (VIBench), specifically designed to assess the system’s performance in non-API-interactable applications across diverse domains such as 3D gaming, recreation, and office scenarios. VIBench complements existing benchmarks by focusing on the agents’ ability to interact with dynamic and visually complex interfaces, thus providing a more strict test of their practical interaction capabilities.
@@ 2 Related work
In this section, we review some related works, with a focus on the domains of LLM-based agents and Autonomous virtual agents.
Figure 2: An overview of MMAC-Copilot, below is the concrete definition of each agent.
@@ 2.1 LLM-based agents
The sunrise field of LLM-based agents, highlighted by projects such as AutoGPT [33] and TaskWeaver [30], marks a significant milestone in the field of LLM research. These agents adopt a human-like approach to problem-solving by decomposing complex tasks into manageable subtasks. The capability demonstrates a profound advancement in planning, observation, and responsive actions, closely mirroring human cognitive processes [38, 40, 45].
@@
The LangChain Agent [36] further extends these capabilities by enabling LLMs to sequence actions, thus expanding the potential for automation across various domains such as customer service, logistics, and content creation. This strategic utilization of LLMs underscores a move towards more sophisticated, context-aware AI systems.
@@
Developments in multi-agent systems, such as AutoGen [43], MetaGPT [17], and AutoAgents [8], introduce a novel dimension to the field by enabling collaboration and competition among agents. This architecture leverages the unique strengths of individual agents, enhancing system-wide efficiency and adaptability in complex scenarios. Such collaborative frameworks are proving essential for large-scale applications in sectors like healthcare [10, 20], finance [14], and urban planning [4].
@@
The versatility and efficacy of these LLM-based agents are highlighted in a range of application domains: In robotics, agents assist in both navigational tasks and complex manipulations, contributing to more autonomous and adaptive robotic systems [12]. Web manipulation and data extraction have seen significant enhancements, enabling better data-driven decision-making and user interaction models [49, 55]. The gaming industry benefits from more dynamic AI opponents and game management systems, which improve user engagement and game design [13, 35]. Automated data analysis tools powered by LLM agents offer sophisticated insights and predictions, facilitating more effective data handling in business and research [39, 53].
2.2 Autonomous virtual agents
@@
The application of LVM systems for interfacing with Graphical User Interfaces (GUIs) in digital applications have been a key area of research. Wonderland [46] harnessed the capabilities of GPT-4V which delineated by Yang et al. [47] to navigate through mobile application interfaces by analyzing screenshots. Concurrently, the AppAgent [48] utilizes GPT-4V to simulate the actions of mobile users, facilitation the execution of tasks in mobile applications through the analysis of device snapshots. Moreover, the MobileAgent incorporates Optical Character Recognition (OCR) technology to enhance the capabilities of of GPT-4V within a similar mobile agent architecture. This strategic enhancement allows MobileAgent to attain task completion rates on par with human operators. In contrast, CogAgent [18] constructs a specifically engineered LVM for interpreting and navigating GUIs. UFO [51] utilizes the windows inspect tool to acquire component and control information from system applications, providing it as context to GPT-4V to fully leverage its planning capabilities.
@@ 3 MMAC-Copilot Framework
12:     end if
13:end for
14:Verify if $S_{i}$ satisfies $R$
15:if not satisfied then
16:     Retry
17:else
18:     Break loop and finalize task
19:end if
@@ 3.1 Specialization of agents
Empirical studies suggest that diversity within human groups contributes to a variety of perspectives, which in turn enhances the group’s performance across various tasks [6, 41, 42]. Also, contemporary research indicates that assigning specific roles to autonomous agents can improve their effectiveness [22, 29, 31]. Thus, in our framework, we have defined six agents, namely: Planner, Librarian, Programmer, Viewer, Video Analyst, and Mentor, as shown in Figure 2. The followings are detailed introductions to each agent:
  • Planner: Employs GPT-4 to strategically manage and allocate tasks among agents, optimizing workflow efficiency. In contrast to other frameworks, our Planner is not required to decompose tasks in to indivisible atomic subtasks. We recognize Planner operating solely in the textual modality may not effectively integrate visual information such as UI elements, potentially leading to hallucination. Therefore, when Planner identifies a subtask which necessitates the incorporation of visual information, it provides only a coarse outline of the subtask. It allows each agent to leverage its own expertise, enhancing the adaptability of the overall framework.
  • Librarian: Utilizes GPT-4 and API for information retrieval, enabling it to answer queries and provide foundational knowledge. It possesses the capability to rapidly process vast amounts of information, thereby supporting user requests and providing related context with other agent.
  • Programmer: Responsible for coding and executing scripts, including Python and Bash commands, to interact with software environments directly. The Programmer’s role is crucial in implementing precise operating systems interactions that adapt to diverse requirements. Key to the Programmer’s efficacy is its ability to refine existing code through a mechanism designed to diagnose and resolve issues effectively. When code errors are identified, Programmer analyzes the source of failure, applies necessary modifications, and maintains the original code structure as much as possible. This refined process includes a detailed explanation of identified issues, enhancing the clarity and functionality of the code without introducing new exception, and improving the handling of existing ones. Furthermore, Programmer can generate accurate Python invocation statements. By deeply understanding the context of the provided Python class, analyzing task descriptions, and interpreting method parameter details to ensure the correct application. The result is a syntactically correct Python statement that calls the methods with appropriate parameters. Lastly, Programmer’s capabilities are complemented by a critical evaluation mechanism. This mechanism assesses the refined and executed code against the user’s specified task requirements. It evaluates the code’s functionality, alignment with the task, and feedback provided by the environment, culminating in a comprehensive reasoning process. The interactions within the Programmer’s mechanism can be mathematically represented as follows:
@@
      $E$ represents the error analysis and feedback.
      $R$ is refinement mechanism.
      $C_{\text{mod}}$ is the modified code.
      $\delta$ includes environmental variables such as the current working directory or system configurations.
      $X$ represents execution function.
      Out includes the runtime output and any runtime errors.
      $T$ is the original task requirements.
      $V$ represents the evaluation mechanism.
      Judge is a boolean indicating if the task is accomplished.
      Score is an integer rating the code’s generality and robustness. If the score is above the specified threshold, the code for this run is retained.
  • Viewer: Integrates GPT-4V for interpreting complex visual information from screenshots. Coupled with SeeClick [9] model, Viewer could translates visual analysis into actionable commands, such as click on identified UI elements, bridging the gap between visual understanding and operating systems interaction.
  • Video Analyst: Uses Gemini Vision to process and analyze video content, extracting critical visual data. It is significant for subtasks which depend on video-based information, enhancing the model’s ability to generalize across different modal. Video Analyst interprets queries, retrieves relevant details from video streams, and delivers accurate responses. These capabilities are crucial for providing contextual insights, assisting other agents in understanding the current statement. In summary, Video Analyst is adept at handling complex scenarios where video data is key.
  • Mentor: Powered by GPT-4V, Mentor provides strategic oversight and troubleshooting support after the execution processes. To be specific, Mentor firstly analyzes screenshots of the current application window to provide detailed description of visible state. And then Mentor identified whether the intended change or commands have successfully impacted the application as expected. Additionally, Mentor generates insights which are derived from a deep analysis of the outcomes and are structured as feedback to be incorporated into the next round.

### v3

# MMAC-Copilot: Multi-modal Agent Collaboration Copilot

@@ Abstract
Large language model agents that interact with PC applications often face limitations due to their singular mode of interaction with real-world environments, leading to restricted versatility and frequent hallucinations. To address this, we propose the Multi-Modal Agent Collaboration framework (MMAC-Copilot), a framework utilizes the collective expertise of diverse agents to enhance interaction ability with application. The framework introduces a team collaboration chain, enabling each participating agent to contribute insights based on their specific domain knowledge, effectively reducing the hallucination associated with knowledge domain gaps. We evaluate MMAC-Copilot using the GAIA benchmark and our newly introduced Visual Interaction Benchmark (VIBench). MMAC-Copilot achieved exceptional performance on GAIA, with an average improvement of 6.8% over existing leading systems. VIBench focuses on non-API-interactable applications across various domains, including 3D gaming, recreation, and office scenarios. It also demonstrated remarkable capability on VIBench. We hope this work can inspire in this field and provide a more comprehensive assessment of Autonomous agents. The anonymous Github is available at Anonymous Github
MMAC-Copilot: Multi-modal Agent Collaboration Copilot
Zirui Song1,2∗Yaohang Li2thanks: Equal contributions.Meng Fang4∗Yanda Li2Zhenhao Chen1 Yuan Huang3Zhecheng Shi3Xiuying Chen1Ling Chen2 1 Mohamed bin Zayed University of Artificial Intelligence 2 University of Technology Sydney 3 Northeastern University 4 University of Liverpool
@@ 1 Introduction
Large Language Models (LLMs) Achiam et al. (2023); Anil et al. (2023); Anthropic (2023); Driess et al. (2023); Ouyang et al. (2022); Touvron et al. (2023a, b) has demonstrated unprecedented capabilities to tackle complex tasks that require human-like reasoning, decision-making, and collaborative efforts Ding et al. (2023); Hao et al. (2023); Park et al. (2023); Qian et al. (2023); Xi et al. (2023). Soon afterwards, Large Vision Models (LVMs) Cai et al. (2023); Li et al. (2024); Liu et al. (2023, 2024); Luo et al. (2023); Zhou et al. (2023a) expand beyond traditional text-based processing by integrating a visual dimension into LLMs. The applications of LVMs have reached diverse domains, with one notable example being their pivotal usage in autonomous agents Wu et al. (2024).
Applying large language models (LLMs) in operating systems faces two primary challenges. First, their ability to acquire real-world environmental information is constrained by reliance on a single text modality, limiting the breadth of applications they can effectively manage. Second, hallucinations stemming from knowledge domain gaps can lead to flawed planning by LLM agents, with potentially severe consequences. Agents typically require both a unified interaction interface and a high level of generalization to adapt to the diverse and complex range of applications within an operating system. While previous works Tan et al. (2024); Wu et al. (2024); Zhang et al. (2024) have leveraged scripting and LLMs’ planning abilities to achieve success in certain scenarios, these challenges persist and must be addressed for broader applicability.
To address the aforementioned challenges, we present MMAC-Copilot, a framework designed to enhance information exchange and collaboration among multi-modal agents. Central to this framework is the team collaboration chain, where agents—each driven by an LLM—adapt the initial plan based on their specific domain expertise. When a user query is received, the Planner formulates an initial plan based on the query’s requirements. The plan is then iteratively refined through collaboration with the other agents, ensuring that insights from each agent’s domain expertise continuously adapt and improve the execution process. As illustrated in Figure 1, MMAC-Copilot integrates five key agents:
  • Planner: Specializes in strategic planning, resource allocation, and ensuring that all agents’ efforts are aligned toward achieving the target.
  • Librarian: Excels in question-answering tasks and API-based information retrieval.
  • Programmer: Handles coding tasks, with expertise in executing Bash commands and Python scripts.
  • Viewer: Focuses on processing visual information, understanding image content, and performing click-based interactions.
  • Video Analyst: Analyzes video content to extract key events.
In addition to these core agents, an auxiliary agent, the Mentor, supervises the system’s interactions with the environment, reviewing the user interface after each interaction to confirm whether the desired outcomes have been achieved.
@@
We evaluate MMAC-Copilot using two benchmarks: GAIA Mialon et al. (2023) and our newly introduced Visual Interaction Benchmark (VIBench). GAIA is designed to assess general artificial intelligence assistants, and MMAC-Copilot achieved scores of 45.16, 20.75, and 6.12 on level-1, level-2, and level-3 tasks (as detailed in Appendix C), outperforming previous methods by 6.8%. VIBench, developed specifically for this study, evaluates performance in non-API-interactable applications across domains such as 3D gaming, recreation, and office environments. The results demonstrate MMAC-Copilot’s exceptional adaptability and precision in handling complex interfaces.
@@
First, we propose MMAC-Copilot, a collaborative framework involving specialized agents designed to navigate and perform general application tasks. MMAC-Copilot enhances autonomous agents’ ability to interact with user interfaces when processing tasks.
Next, to mitigate the incorrect planning issue caused by LLMs’ hallucination, we introduce the team collaboration chain which allows participating agents to adjust the initial plan based on their domain expertise into MMAC-Copilot. It has demonstrated adaptability in handling a variety of case.
Finally, we develop the Visual Interaction Benchmark (VIBench), specifically designed to assess the system’s performance in non-API-interactable applications across diverse domains such as 3D gaming, recreation, and office scenarios. VIBench complements existing benchmarks by focusing on the agents’ ability to interact with visually complex user interfaces.
@@ 2 Related work
In this section, we review some related works, with a focus on the domains of LLM-based agents and autonomous agents.
Figure 1: An overview of MMAC-Copilot framework.
@@ 2.1 LLM-based agents
The sunrise field of LLM-based agents, highlighted by projects such as AutoGPT Significant-Gravitas (2023) and TaskWeaver Qiao et al. (2023), marks a significant milestone in the field of LLM research. These agents adopt a human-like approach to problem-solving by decomposing complex tasks into manageable subtasks. The capability demonstrates a profound advancement in planning, observation, and responsive actions, closely mirroring human cognitive processes Touvron et al. (2023b); Wang et al. (2024); Xi et al. (2023).
@@
The LangChain Agent Topsakal and Akinci (2023) further extends these capabilities by enabling LLMs to sequence actions, thus expanding the potential for automation. This strategic utilization of LLMs underscores a move towards more sophisticated, context-aware AI systems.
@@
Developments in multi-agent systems, such as AutoGen Wu et al. (2023), MetaGPT Hong et al. (2023a), and AutoAgents Chen et al. (2023), introduce a novel dimension to the field by enabling collaboration and competition among agents. This architecture leverages the unique strengths of individual agents, enhancing system-wide efficiency and adaptability in complex scenarios. Such collaborative frameworks are proving essential for large-scale applications in sectors like healthcare Dhatterwal et al. (2023); Jones et al. (2022), finance Gautier et al. (2023), and urban planning Bae et al. (2022).
@@
The versatility and efficacy of these LLM-based agents are highlighted in a range of application domains: In robotics, agents assist in both navigational tasks and complex manipulations, contributing to more autonomous and adaptive robotic systems Driess et al. (2023). Web manipulation and data extraction have seen significant enhancements, enabling better data-driven decision-making and user interaction models Yao et al. (2022); Zhou et al. (2023b). The gaming industry benefits from more dynamic AI opponents and game management systems, which improve user engagement and game design Fan et al. (2022); Tan et al. (2024). Automated data analysis tools powered by LLM agents offer sophisticated insights and predictions, facilitating more effective data handling in business and research Wang et al. (2023); Zhang et al. (2023b).
2.2 Autonomous agents
@@
The application of LVM systems for interfacing with Graphical User Interfaces (GUIs) in digital applications have been a key area of research. Wonderland Yan et al. (2023) harnessed the capabilities of GPT-4V which delineated by Yang et al. Yang et al. (2023b) to navigate through mobile application interfaces by analyzing screenshots. Concurrently, the AppAgent Yang et al. (2023a) utilizes GPT-4V to simulate the actions of mobile users, facilitation the execution of tasks in mobile applications through the analysis of device snapshots. Moreover, the MobileAgent incorporates Optical Character Recognition (OCR) technology to enhance the capabilities of of GPT-4V within a similar mobile agent architecture. This strategic enhancement allows MobileAgent to attain task completion rates on par with human operators. In contrast, CogAgent Hong et al. (2023b) constructs a specifically engineered LVM for interpreting and navigating GUIs. UFO Zhang et al. (2024) utilizes the windows inspect tool to acquire component and control information from system applications, providing it as context to GPT-4V to fully leverage its planning capabilities.
@@ 3 MMAC-Copilot Framework
12:     else if $F_{i}$ indicates no adjustments are needed then
13:         if $s$ is the last subtask in $P_{0}$ then
14:              Terminate loop
15:         else
16:              Continue to the next subtask
17:         end if
18:     end if
19:end for
@@ 3.1 Specialization of agents
Empirical studies suggest that diversity within human groups contributes to a variety of perspectives, which in turn enhances the group’s performance across various tasks Bransford and Stein (1993); Williams and O’Reilly III (1998); Woolley et al. (2015). Also, contemporary research indicates that assigning specific roles to autonomous agents can improve their effectiveness Li et al. (2023); Qian et al. (2023); Salewski et al. (2024). Thus, in our framework, we have defined six agents, namely: Planner, Librarian, Programmer, Viewer, Video Analyst, and Mentor, as shown in Figure 1. The followings are detailed introductions to each agent:
Planner: Employs LLM to strategically manage and allocate tasks among agents, optimizing workflow efficiency. In contrast to other frameworks, our Planner is not required to decompose tasks into indivisible atomic subtasks. We recognize Planner operating solely in the textual modality may not effectively integrate visual information such as UI elements, potentially leading to hallucination. Therefore, when Planner identifies a subtask which necessitates the incorporation of visual information, it provides only a coarse outline of the subtask. It allows each agent to leverage its own expertise, enhancing the adaptability of the overall framework.
Librarian: Utilizes LLM and API for information retrieval, enabling it to answer queries and provide foundational knowledge. It possesses the capability to rapidly process vast amounts of information, thereby supporting user requests and providing related context with other agent.
Programmer: Be responsible for writing and executing scripts, such as Python and Bash, to directly interact with software environments. This role is essential for tailoring applications to meet diverse requirements. The Programmer refines existing code by diagnosing and resolving issues. This process involves clarifying identified problems, enhancing code functionality, handling exceptions better, and providing detailed explanations.
Accurate Python statements are generated by understanding the context of Python classes and analyzing method parameters, ensuring syntactically correct calls. Also a evaluation mechanism further complements the Programmer’s capabilities, assessing the refined code against task requirements, its functionality, and alignment, leading to a comprehensive review process. More detailed Prompt has been listed on Appendix B.
The interactions within the Programmer’s mechanism are mathematically represented as:
@@
  • $E$ stands for error analysis.
  • $R$ is the refinement process.
  • $C_{\text{mod}}$ represents the modified code.
  • $\delta$ includes environmental variables.
  • $X$ is the execution function.
  • Out includes runtime outputs and errors.
  • $T$ corresponds to the original task requirements.
  • $V$ is the evaluation process.
  • Judge determines if the task is accomplished.
  • Score assesses the code’s generality and robustness, with higher scores indicating retention.
Viewer: The Viewer module integrates GPT-4V for the sophisticated interpretation of complex visual data extracted from screenshots. By coupling this capability with the SeeClick model Cheng et al. (2024), the Viewer effectively translates visual analyses into executable commands—such as initiating clicks on identified user interface elements. This integration bridges the gap between visual comprehension and application interaction, enabling automated manipulation of graphical interfaces based on visual understanding.
Video Analyst: The Video Analyst leverages the Gemini Vision to process and analyze video content, extracting critical visual information relevant to the task domain. This module is essential for subtasks that depend on video-based data, enhancing the model’s ability to generalize across different modalities. By interpreting queries and retrieving pertinent details from video streams, the Video Analyst delivers precise and contextually appropriate responses. These capabilities are crucial for providing comprehensive contextual insights and assisting other agents in understanding the current operational state. Consequently, the Video Analyst adeptly handles complex scenarios where video data is paramount.
Mentor: Powered by GPT-4V, the Mentor agent plays a pivotal role in overseeing the system’s execution processes, providing strategic oversight and troubleshooting support. After each subtask is executed, the Mentor analyzes screenshots from the current application window, offering a detailed description of the visible state. It evaluates whether the intended changes or commands have successfully impacted the application as expected. If the analysis indicates that adjustments are needed, the Mentor generates feedback based on these observations, which is then incorporated into the system’s planning process for the next iteration.
This feedback loop allows the planner to revise the current plan and reattempt the adjusted task. However, if no adjustments are necessary, the system proceeds to the next subtask in the sequence. In the case where the last subtask is reached and no further changes are required, the system terminates the process. This iterative mechanism of feedback and improvement not only ensures the successful completion of tasks but also enhances the overall system performance by continuously refining actions based on empirical data.